OpenStudy (anonymous):

Complete the square to solve the equation. x^2+2x-14=10 a. x=5; x=4 b. x=1 ; x=4 c. x=1; x=3 d. x=-6; x=4

OpenStudy (lalaly):

(x+1)^2-1-14=10 (x+1)^2=25 x+1=5 , x+1=-5 x=4 x=-6 so its (d)

OpenStudy (anonymous):

Thanks!

OpenStudy (lalaly):

:D :D